# Vital Signs Monitoring with mmWave OFDM JCAS System

### v1

@@ Abstract
Index Terms:
Vital signs monitoring, millimeter-wave radar, OFDM, heart rate detection, breathing measurement, contactless sensing, joint communication and sensing
@@ I-A Principles of Contactless Respiration Rate and Heart Rate Monitoring
The detected signal is, for the most part, reflected from the skin of the person and not from the internal organs. It results from the high attenuation of EM waves by body tissues. Additionally, for radio waves below 18 GHz, the magnitude of the reflection coefficient $\Gamma_{\text{air/skin}}$ between air and skin is above 0.7, which means that the air-skin interface reflects more than 70% of the energy [1]. In the case of millimeter waves (mmWaves), less energy is reflected, but these waves are predominantly absorbed in the first few centimeters of the tissue. Therefore, even though the movements of the heart are greater than those of the skin, the amount of signal reflected from the heart walls is so low that it has little impact on the received signal [1].
@@
Apart from the measurement based on phase variation, an alternative approach for vital signs monitoring involves tracking changes in the magnitude of the received signal caused by vital signs. However, magnitude-based measurements are generally considered less robust because the small movements involved cause only minor variations in signal power, which are difficult to measure [2]. Measuring HR using signal strength change can be practically challenging, but it is more feasible for applications such as tracking the number of people in a room.
@@ II-A Hospital and Post-ICU Care
In hospital settings, studies have demonstrated that continuous vital signs monitoring using wearable wireless devices can lead to improved patient outcomes. Wireless monitoring has been associated with shorter hospital stays, fewer nurse-to-physician calls in surgical patients, and a reduction in rapid response team calls, suggesting improved early recognition and management of clinical decline [3].
@@
Wireless vital signs monitoring provides a valuable safety net during this vulnerable post-ICU transition phase. Continuous tracking of key parameters can detect subtle changes that precede clinical deterioration. Several studies support this approach, showing that continuous monitoring in step-down units can significantly reduce ‘failure-to-rescue’ events and unplanned ICU readmissions [3].
Moreover, the use of mobile or wearable wireless systems facilitates early discharge planning. It enables step-down care units or even home-based telemonitoring programs to safely monitor cardiac patients while reducing hospital stay durations and overall costs. This strategy aligns with modern heart failure and post-acute care guidelines that emphasize early mobilization and structured follow-up [4].
@@ II-B Long-Term Care and Home Monitoring
In long-term care facilities and home settings, wireless monitoring systems offer a way to maintain continuous oversight of elderly patients’ health, reducing the need for frequent in-person visits and enabling prompt response to early warning signs. A recent study highlighted the potential of wearable wireless sensors in out-of-hospital settings, noting that such devices can signal when a person should seek additional care, thereby improving outcomes for seniors living independently. Personalized health-monitoring systems that combine rule-based logic with case-based reasoning may also enhance tailored care and early intervention [5].
Devices for continuous vital sign monitoring may benefit various groups. In hospital inpatients, this allows for accurate and remote monitoring while minimizing disruptions and reducing the need for manual measurements. For infectious disease patients, they help limit exposure risk for healthcare workers. Furthermore, they increase patient mobility and comfort by reducing the number of wires and sensors needed [6].
@@ II-C Future Directions in Healthcare
A promising future direction for wireless vital signs monitoring is the development of systems that do not require the patient to wear any device. Removing the need for physical sensors would enhance comfort and ease of use, particularly for the elderly or individuals with cognitive impairments. It would also address common challenges such as forgetting to wear the device, accidental dislodgement, or the need for regular charging and maintenance. This contactless approach could simplify monitoring workflows, reduce caregiver burden, and provide a more seamless and less intrusive experience for patients—particularly in long-term care and home settings [4, 5]. With the rapid technological improvement and growing demand for convenient alternatives to existing contact-based health monitoring systems, the U.S. market for vital signs monitoring devices is projected to exceed USD 11 billion by 2028 [7].
@@ III Related Work
One of the most promising technologies for contactless measurement of human VS is the frequency-modulated continuous wave (FMCW) radar. FMCW radars are commercially available and widely utilized, especially in the automotive industry, as part of Advanced Driver-Assistance Systems (ADAS). The non-linear components of the system, namely power amplifiers and low-noise amplifiers, operate in their linear region in FMCW radars. This is a result of the fact that the FMCW signal has a constant envelope with a peak-to-average power ratio (PAPR) of 0 dB [9]. In the literature, different frequencies have been tested, including 9.6 GHz, 24 GHz, and 77 GHz, with the 77–81 GHz band being the most widely used for this purpose. FMCW radars can estimate the distance to a target with the range resolution of $\Delta R=c/2B$, where $c$ is the speed of light in vacuum, and $B$ is the bandwidth in Hz. This imposes high bandwidth requirements, especially in the context of VS monitoring applications, where body movement resulting from the heartbeat and breathing is on the order of centimeters or even millimeters.
@@
The article [9] presents the results of experiments conducted using a Texas Instruments mm-wave FMCW radar (AWR1443) operating in the 77–81 GHz frequency band. The experiment utilizes a single transmitter-receiver (Tx/Rx) pair with a chirp duration of 64 µs and a repetition period of $T_{c}=50$ ms. The slope of each chirp is 70 MHz/µs with a bandwidth of 3.99 GHz. The maximum unambiguous distance between the patient and the radar is 2.14 m, which means that for greater distances, the radar will report values modulo 2.14 m. The measurements of BR and HR are taken after gathering samples from 256 chirp sequences, which gives the observation duration of 12.8 s. For optimization, the observation window utilizes 128 samples from the previous observation, resulting in new BR and HR measurements being effectively produced every 6.4 seconds. The study examines the performance of the radar in a scenario where the patient lies on their bed, facing up at the radar, which is mounted above. A wearable device is used as a reference for the tested device. The FMCW radar achieved 80% accuracy for HR estimation and 94% accuracy for breathing rate at a distance of 1.7 m with an output power of 15.8 mW. The results are comparable with those of others, as article [10] reports achieving accuracies of 87.2% and 91.08% in the estimation of HR and BR, respectively. The setup included ultra-wideband 80 GHz FMCW with 10 GHz of bandwidth.
@@
An interesting approach for differentiating between multiple closely spaced subjects is presented in  [11]. The available bandwidth of the 24 GHz FMCW radar used in the study was 250 MHz, giving a theoretical resolution of 60 cm. With the help of the MUSIC algorithm for signal separation, the proposed method can detect the vital signs of two people located just 40 cm apart (in terms of their distance from the radar), which exceeds the theoretical range resolution limit. When subjects were closer than 40 cm, they could be spatially identified, but the vital sign signals interfered, and the results were not accurate. Larger non-vital movements posed a challenge to the radar, and the authors suggested combining MUSIC with micro-Doppler methods for extracting only VS motions.
@@
A more recent method for measuring VS in a multi-target scenario is presented in [12]. The authors utilize a 77 GHz FMCW radar with a 3.5 GHz bandwidth. The authors used Moving Target Indication for stationary clutter suppression, Constant False Alarm Rate Detection for detecting the distance to the target peaks, and K-means clustering for grouping detected peaks into clusters representing people. The developed method was tested in an office scenario with eight persons (four scenarios of two persons each), with distances from one to two meters. It achieved mean absolute errors of 1.22 bpm for BR and 3.62 bpm for HR, and mean accuracies of 94.2% and 95.1%, respectively.
@@
A notable advancement in Wi-Fi sensing is presented in [13]. The authors examine the use of modified Wi-Fi devices for simultaneously measuring the BR of two individuals. The scenario involves using an antenna array at the receiver, enabling the resolution of two targets. The carrier frequency is 5.32 GHz, and the receiver features four microstrip Yagi antennas. The setup includes an OFDM communication system. An algorithm evaluates the variance of the channel frequency response for each subcarrier. The subcarriers with the highest variance are most sensitive to motion, so only they are used for sensing. The obtained measurements are highly accurate, particularly when the device is located close to the subjects. For distances below 1.5 meters, the system achieves accuracy above 95% in both single-person and two-person scenarios. Accuracy degrades with distance, more rapidly in the two-target case.
Another significant contribution is presented in [14], where the authors analyzed human BR and HR during sleep using off-the-shelf Wi-Fi. The setup included an 802.11n commercial TP-Link TL-WDR4300 access point transmitting the signal. The system used CSI from 30 subcarriers evenly distributed across the 56 subcarriers of a 20 MHz OFDM channel. For BR estimation, the proposed algorithm achieved absolute errors of less than 0.4 bpm in over 90% of measurements for AP–device distances ranging from 2 to 7 meters. The system achieved HR estimation errors of less than 4 bpm in more than 90% of cases.
In [15], the authors described Wital, a low-cost vital signs monitoring system based on the same commercial Wi-Fi hardware as described in [14]. The system used CSI data for VS measurement in a non-line-of-sight scenario. A scenario was set up by placing a lead sheet between the transmitter and the receiver. The system achieved a mean BR estimation error of 0.498 bpm with 96.89% accuracy, and a mean HR estimation error of 3.531 bpm with 94.71% accuracy.
@@ IV Measurement Setup
The measurements were taken in the laboratory of the Fraunhofer Institute for Integrated Circuits IIS in Ilmenau, using a software-defined radio (SDR) platform built and maintained by researchers from the hosting institute and the Institute of Information Technology at Technische Universität Ilmenau. The system’s architecture is described in [16]. The setup consisted of a bistatic, single-input single-output (SISO) system with two horn antennas, one transmitter (Tx) and one receiver (Rx), specifically, the TACTRON HA-QR-18000-40000-2x2.92F (18–40 GHz, 13 dBi gain at 26 GHz). The diagram of the laboratory setup is presented in Fig.  1. Fig.  2 shows the Tx and Rx antennas used in the setup. The waveform was a custom OFDM signal with 1024 subcarriers and a bandwidth of 1 GHz. Its frequency-domain representation is shown in Fig. 3. The subcarriers did not carry any data. Instead, their phases were specifically adjusted to reduce the peak-to-average power ratio, optimizing them for signal processing purposes. The system operated at a carrier frequency of 26.5 GHz, corresponding to a wavelength of approximately 1.13 cm, which is comparable to the amplitudes of body movements caused by HR and respiration. The bandwidth determines the range resolution, which is equal to 15 cm. Therefore, measuring vital signs would not be possible by tracking the changing distance between the antenna and the human body using the time of flight of the signal. The range bin is 7.5 cm, which is too large to capture breathing and heart-induced movements. Each OFDM symbol was sampled 2500 times. The system recorded the reflected signal for multiple OFDM pulses. The recorded data was averaged in slow time because the raw data rate was nearly 20 GB/s. The averaging factor of 100 in the measurements resulted in a lower data rate, allowing for longer continuous patient monitoring. The most essential parameters of the devices used are summarized in Table I.
@@ IV-A Measurement Scenarios
A set of measurement scenarios was designed to evaluate the radar’s performance in different conditions. These scenarios illustrate the potential applications for vital signs monitoring in hospital conditions. Indoor stationary and dynamic scenarios were considered. The experiments aimed to verify the influence of various factors, including the angle and distance between the patient and the radar, the person’s position and activity, as well as line-of-sight conditions. Vital signs were also measured using reference devices to provide ground truth for the radar performance. Holter EKG EDAN SE-2003/201 [17] and a chest-strap Polar H9 Heart Rate Monitor [18] were used for pulse monitoring. Breathing rate was measured by extracting the respiratory motion from the patient’s video.
@@ V Signal Processing
Human vital signs, such as breathing and heartbeat, result in periodic movements of the human body, primarily in the thoracic region, neck, and face. During respiration, the expansion and contraction of the lungs cause the chest wall to move. Heart activity generates periodic changes in blood pressure as it ejects blood into the arterial system, creating a pressure wave propagating through the arterial walls. This pressure change causes the pulsation of arteries as they expand and contract rhythmically. These periodic expansions result in a small mechanical displacement of the surrounding tissue and skin. In areas where the arteries are located just under the skin, their movement can induce vibrations on the surface with an amplitude range of 0.2–0.5 mm and a frequency range of 1–1.34 Hz (approximately 60–80 beats per minute) [19]. The chest surface typically moves within a range of 4–12 mm with a frequency of 0.2–0.34 Hz (approximately 12–20 breaths per minute) [19].
In the experiments, a custom-built radar operating at a center frequency of 26.5 GHz was used for data collection. This places the system in the lower part of the mmWave band, where human skin shows significant reflectivity due to its high water content and dielectric properties. When mmWave electromagnetic radiation contacts the person’s skin, some of the radiation is reflected, and the rest is absorbed near the skin surface [20]. In the lower regions of the mmWave band, for normal incidence—where the wavefront is perpendicular to the skin surface—approximately 40–50% of the incident electromagnetic wave is reflected at the air–skin interface [21]. The actual value depends on the angle of incidence, wave polarization, exact frequency, and water content of the skin.

### v2

@@ I-A Principles of Contactless Respiration Rate and Heart Rate Monitoring
The detected signal is, for the most part, reflected from the skin of the person and not from the internal organs. It results from the high attenuation of EM waves by body tissues. Additionally, for radio waves below 18 GHz, the magnitude of the reflection coefficient $\Gamma_{\text{air/skin}}$ between air and skin is above 0.7, which means that the air-skin interface reflects more than 70% of the energy [Aardal2013]. In the case of millimeter waves (mmWaves), less energy is reflected, but these waves are predominantly absorbed in the first few centimeters of the tissue. Therefore, even though the movements of the heart are greater than those of the skin, the amount of signal reflected from the heart walls is so low that it has little impact on the received signal [Aardal2013].
@@
Apart from the measurement based on phase variation, an alternative approach for vital signs monitoring involves tracking changes in the magnitude of the received signal caused by vital signs. However, magnitude-based measurements are generally considered less robust because the small movements involved cause only minor variations in signal power, which are difficult to measure [Alizadeh2019]. Measuring HR using signal strength change can be practically challenging, but it is more feasible for applications such as tracking the number of people in a room.
@@ II-A Hospital and Post-ICU Care
In hospital settings, studies have demonstrated that continuous vital signs monitoring using wearable wireless devices can lead to improved patient outcomes. Wireless monitoring has been associated with shorter hospital stays, fewer nurse-to-physician calls in surgical patients, and a reduction in rapid response team calls, suggesting improved early recognition and management of clinical decline [Leenen2024].
@@
Wireless vital signs monitoring provides a valuable safety net during this vulnerable post-ICU transition phase. Continuous tracking of key parameters can detect subtle changes that precede clinical deterioration. Several studies support this approach, showing that continuous monitoring in step-down units can significantly reduce ‘failure-to-rescue’ events and unplanned ICU readmissions [Leenen2024].
Moreover, the use of mobile or wearable wireless systems facilitates early discharge planning. It enables step-down care units or even home-based telemonitoring programs to safely monitor cardiac patients while reducing hospital stay durations and overall costs. This strategy aligns with modern heart failure and post-acute care guidelines that emphasize early mobilization and structured follow-up [Clark2023].
@@ II-B Long-Term Care and Home Monitoring
In long-term care facilities and home settings, wireless monitoring systems offer a way to maintain continuous oversight of elderly patients’ health, reducing the need for frequent in-person visits and enabling prompt response to early warning signs. A recent study highlighted the potential of wearable wireless sensors in out-of-hospital settings, noting that such devices can signal when a person should seek additional care, thereby improving outcomes for seniors living independently. Personalized health-monitoring systems that combine rule-based logic with case-based reasoning may also enhance tailored care and early intervention [Ahmed2015].
Devices for continuous vital sign monitoring may benefit various groups. In hospital inpatients, this allows for accurate and remote monitoring while minimizing disruptions and reducing the need for manual measurements. For infectious disease patients, they help limit exposure risk for healthcare workers. Furthermore, they increase patient mobility and comfort by reducing the number of wires and sensors needed [Weller2018].
@@ II-C Future Directions in Healthcare
A promising future direction for wireless vital signs monitoring is the development of systems that do not require the patient to wear any device. Removing the need for physical sensors would enhance comfort and ease of use, particularly for the elderly or individuals with cognitive impairments. It would also address common challenges such as forgetting to wear the device, accidental dislodgement, or the need for regular charging and maintenance. This contactless approach could simplify monitoring workflows, reduce caregiver burden, and provide a more seamless and less intrusive experience for patients—particularly in long-term care and home settings [Clark2023, Ahmed2015]. With the rapid technological improvement and growing demand for convenient alternatives to existing contact-based health monitoring systems, the U.S. market for vital signs monitoring devices is projected to exceed USD 11 billion by 2028 [ResearchMarkets2023].
@@ III Related Work
One of the most promising technologies for contactless measurement of human VS is the frequency-modulated continuous wave (FMCW) radar. FMCW radars are commercially available and widely utilized, especially in the automotive industry, as part of Advanced Driver-Assistance Systems (ADAS). The non-linear components of the system, namely power amplifiers and low-noise amplifiers, operate in their linear region in FMCW radars. This is a result of the fact that the FMCW signal has a constant envelope with a peak-to-average power ratio (PAPR) of 0 dB [alizadeh2019remote]. In the literature, different frequencies have been tested, including 9.6 GHz, 24 GHz, and 77 GHz, with the 77–81 GHz band being the most widely used for this purpose. FMCW radars can estimate the distance to a target with the range resolution of $\Delta R=c/2B$, where $c$ is the speed of light in vacuum, and $B$ is the bandwidth in Hz. This imposes high bandwidth requirements, especially in the context of VS monitoring applications, where body movement resulting from the heartbeat and breathing is on the order of centimeters or even millimeters.
@@
The article [alizadeh2019remote] presents the results of experiments conducted using a Texas Instruments mm-wave FMCW radar (AWR1443) operating in the 77–81 GHz frequency band. The experiment utilizes a single transmitter-receiver (Tx/Rx) pair with a chirp duration of 64 µs and a repetition period of $T_{c}=50$ ms. The slope of each chirp is 70 MHz/µs with a bandwidth of 3.99 GHz. The maximum unambiguous distance between the patient and the radar is 2.14 m, which means that for greater distances, the radar will report values modulo 2.14 m. The measurements of BR and HR are taken after gathering samples from 256 chirp sequences, which gives the observation duration of 12.8 s. For optimization, the observation window utilizes 128 samples from the previous observation, resulting in new BR and HR measurements being effectively produced every 6.4 seconds. The study examines the performance of the radar in a scenario where the patient lies on their bed, facing up at the radar, which is mounted above. A wearable device is used as a reference for the tested device. The FMCW radar achieved 80% accuracy for HR estimation and 94% accuracy for breathing rate at a distance of 1.7 m with an output power of 15.8 mW. The results are comparable with those of others, as article [wang2015novel] reports achieving accuracies of 87.2% and 91.08% in the estimation of HR and BR, respectively. The setup included ultra-wideband 80 GHz FMCW with 10 GHz of bandwidth.
@@
An interesting approach for differentiating between multiple closely spaced subjects is presented in  [lee2019novel]. The available bandwidth of the 24 GHz FMCW radar used in the study was 250 MHz, giving a theoretical resolution of 60 cm. With the help of the MUSIC algorithm for signal separation, the proposed method can detect the vital signs of two people located just 40 cm apart (in terms of their distance from the radar), which exceeds the theoretical range resolution limit. When subjects were closer than 40 cm, they could be spatially identified, but the vital sign signals interfered, and the results were not accurate. Larger non-vital movements posed a challenge to the radar, and the authors suggested combining MUSIC with micro-Doppler methods for extracting only VS motions.
@@
A more recent method for measuring VS in a multi-target scenario is presented in [XUE2023113715]. The authors utilize a 77 GHz FMCW radar with a 3.5 GHz bandwidth. The authors used Moving Target Indication for stationary clutter suppression, Constant False Alarm Rate Detection for detecting the distance to the target peaks, and K-means clustering for grouping detected peaks into clusters representing people. The developed method was tested in an office scenario with eight persons (four scenarios of two persons each), with distances from one to two meters. It achieved mean absolute errors of 1.22 bpm for BR and 3.62 bpm for HR, and mean accuracies of 94.2% and 95.1%, respectively.
@@
A notable advancement in Wi-Fi sensing is presented in [Guan2023]. The authors examine the use of modified Wi-Fi devices for simultaneously measuring the BR of two individuals. The scenario involves using an antenna array at the receiver, enabling the resolution of two targets. The carrier frequency is 5.32 GHz, and the receiver features four microstrip Yagi antennas. The setup includes an OFDM communication system. An algorithm evaluates the variance of the channel frequency response for each subcarrier. The subcarriers with the highest variance are most sensitive to motion, so only they are used for sensing. The obtained measurements are highly accurate, particularly when the device is located close to the subjects. For distances below 1.5 meters, the system achieves accuracy above 95% in both single-person and two-person scenarios. Accuracy degrades with distance, more rapidly in the two-target case.
Another significant contribution is presented in [wifi-posture-sleep], where the authors analyzed human BR and HR during sleep using off-the-shelf Wi-Fi. The setup included an 802.11n commercial TP-Link TL-WDR4300 access point transmitting the signal. The system used CSI from 30 subcarriers evenly distributed across the 56 subcarriers of a 20 MHz OFDM channel. For BR estimation, the proposed algorithm achieved absolute errors of less than 0.4 bpm in over 90% of measurements for AP–device distances ranging from 2 to 7 meters. The system achieved HR estimation errors of less than 4 bpm in more than 90% of cases.
In [gu2021realtime], the authors described Wital, a low-cost vital signs monitoring system based on the same commercial Wi-Fi hardware as described in [wifi-posture-sleep]. The system used CSI data for VS measurement in a non-line-of-sight scenario. A scenario was set up by placing a lead sheet between the transmitter and the receiver. The system achieved a mean BR estimation error of 0.498 bpm with 96.89% accuracy, and a mean HR estimation error of 3.531 bpm with 94.71% accuracy.
@@ IV Measurement Setup
The measurements were taken in the laboratory of the Fraunhofer Institute for Integrated Circuits IIS in Ilmenau, using a software-defined radio (SDR) platform built and maintained by researchers from the hosting institute and the Institute of Information Technology at Technische Universität Ilmenau. The system’s architecture is described in [engelhardt2024sdr]. The setup consisted of a bistatic, single-input single-output (SISO) system with two horn antennas, one transmitter (Tx) and one receiver (Rx), specifically, the TACTRON HA-QR-18000-40000-2x2.92F (18–40 GHz, 13 dBi gain at 26 GHz). The diagram of the laboratory setup is presented in Fig.  1. Fig.  2 shows the Tx and Rx antennas used in the setup. The waveform was a custom OFDM signal with 1024 subcarriers and a bandwidth of 1 GHz. Its frequency-domain representation is shown in Fig. 3. The subcarriers did not carry any data. Instead, their phases were specifically adjusted to reduce the peak-to-average power ratio, optimizing them for signal processing purposes. The system operated at a carrier frequency of 26.5 GHz, corresponding to a wavelength of approximately 1.13 cm, which is comparable to the amplitudes of body movements caused by HR and respiration. The bandwidth determines the range resolution, which is equal to 15 cm. Therefore, measuring vital signs would not be possible by tracking the changing distance between the antenna and the human body using the time of flight of the signal. The range bin is 7.5 cm, which is too large to capture breathing and heart-induced movements. Each OFDM symbol was sampled 2500 times. The system recorded the reflected signal for multiple OFDM pulses. The recorded data was averaged in slow time because the raw data rate was nearly 20 GB/s. The averaging factor of 100 in the measurements resulted in a lower data rate, allowing for longer continuous patient monitoring. The most essential parameters of the devices used are summarized in Table I.
@@ IV-A Measurement Scenarios
A set of measurement scenarios was designed to evaluate the radar’s performance in different conditions. These scenarios illustrate the potential applications for vital signs monitoring in hospital conditions. Indoor stationary and dynamic scenarios were considered. The experiments aimed to verify the influence of various factors, including the angle and distance between the patient and the radar, the person’s position and activity, as well as line-of-sight conditions. Vital signs were also measured using reference devices to provide ground truth for the radar performance. Holter EKG EDAN SE-2003/201 [edan-holter] and a chest-strap Polar H9 Heart Rate Monitor [polar-h9] were used for pulse monitoring. Breathing rate was measured by extracting the respiratory motion from the patient’s video.
@@ V Signal Processing
Human vital signs, such as breathing and heartbeat, result in periodic movements of the human body, primarily in the thoracic region, neck, and face. During respiration, the expansion and contraction of the lungs cause the chest wall to move. Heart activity generates periodic changes in blood pressure as it ejects blood into the arterial system, creating a pressure wave propagating through the arterial walls. This pressure change causes the pulsation of arteries as they expand and contract rhythmically. These periodic expansions result in a small mechanical displacement of the surrounding tissue and skin. In areas where the arteries are located just under the skin, their movement can induce vibrations on the surface with an amplitude range of 0.2–0.5 mm and a frequency range of 1–1.34 Hz (approximately 60–80 beats per minute) [Shafiq2014]. The chest surface typically moves within a range of 4–12 mm with a frequency of 0.2–0.34 Hz (approximately 12–20 breaths per minute) [Shafiq2014].
In the experiments, a custom-built radar operating at a center frequency of 26.5 GHz was used for data collection. This places the system in the lower part of the mmWave band, where human skin shows significant reflectivity due to its high water content and dielectric properties. When mmWave electromagnetic radiation contacts the person’s skin, some of the radiation is reflected, and the rest is absorbed near the skin surface [Owda2017]. In the lower regions of the mmWave band, for normal incidence—where the wavefront is perpendicular to the skin surface—approximately 40–50% of the incident electromagnetic wave is reflected at the air–skin interface [Wu2015]. The actual value depends on the angle of incidence, wave polarization, exact frequency, and water content of the skin.